# Present address: Physics Dept., U.C. Santa Barbara, Santa Barbara, California 93106, USA and Google Quantum AI, 301 Mentor Dr, Goleta, CA 93111, USA

### v2

@@ Abstract
Notably, the suppression of the burst detection rate by gap engineering is only about a factor of five. This observation is inconsistent with the QP energy thermalizing with the mixing chamber of the fridge immediately after the impact. If QPs were thermalized – as for resident QPs in Fig. 1 – $\>\delta\Delta$ would suppress the burst detection rate by about a factor of $\exp(\left[\>\delta\Delta-hf_{q}\right]/\>k_{\mathrm{B}}T)\sim 10^{4}$ in the big-$\>\delta\Delta$ device. Here, $T=25\>\mathrm{mK}$ is the base temperature of our fridge.
@@
Next, we consider how $T_{1}$ of the qubit returns to its steady state after the impact. To this end, we use continuous monitoring data from the measurement of Fig. 2(b-d). We identify the bursts using the threshold shown in Fig. 2(c). Averaging the measurement outcomes between many observed burst events allows us to determine the qubit relaxation rate $\Gamma_{10}$ as a function of time after the impact [noauthor_see_nodate]. Fig. 2(e) shows the results. For both devices, the relaxation rate returns to the steady state rougly exponentially. The associated time constants $\sim 0.7\>\mathrm{ms}$ are insensitive to the gap engineering.
In addition to the qubit relaxation rate $\Gamma_{10}(t)$ after the impact, we similarly measure the qubit excitation rate $\Gamma_{01}(t)$. To this end, we repeat the measurement similar to that described in Fig. 2(b), except without resetting the qubit to |$\ket{1}$. From the two rates, one may extract the time dependence of the QP density, $\>x_{\mathrm{QP}}(t)$, and the qubit temperature $T_{q}(t)$, see the inset of Fig. 2(e) for the latter quantity. Here, the qubit temperature is defined as $\>k_{\mathrm{B}}T_{q}=hf_{q}/\mathrm{ln}\left(\Gamma_{10}/\Gamma_{01}\right)$. The time dependence of $\>x_{\mathrm{QP}}$ is similar to that of $\Gamma_{10}(t)$. In about a millisecond, the QP density decays to a level where QPs no longer limit the qubit $T_{1}$.
@@
The qubit temperature $T_{q}$ surges to about 90 mK after the impact, significantly exceeding the fridge’s base temperature. The qubit temperature returns to its steady state exponentially, with a time constant of about $6\>\mathrm{ms}$. This time scale is almost 10 times longer than the $0.7\>\mathrm{ms}$ time constant for the decay of $\>x_{\mathrm{QP}}$. The temperature thus remains elevated even long after the generated by the impact non-equilibrium QPs are gone. This points to a dominant role of the substrate properties in determining the evolution of $T_{q}(t)$. Indeed, the observed increase in qubit temperature to $90\>\mathrm{mK}$ is consistent with the substrate phonon heat capacity for a typical radiation impact energy of $\sim 100\mathrm{keV}-1\>\mathrm{MeV}$ [fowler_spectroscopic_2024, noauthor_see_nodate]. The decay constant for the qubit temperature is determined by the rate of phonon escape from the system, likely through the chip clamps.
@@
The elevated substrate temperature after burst explains the weak sensitivity of burst occurrence to the gap difference. As previously described, the suppression of QP-induced qubit relaxation by $\>\delta\Delta$ is controlled by the activation exponent $\exp(-[\delta\Delta-hf_{q}]/\>k_{\mathrm{B}}T)$. For the gap difference to be effective, $\delta\Delta-hf_{q}$ must be much larger than $\>k_{\mathrm{B}}T$. For our big-$\delta\Delta$ device, $(\delta\Delta-hf_{q})/\>k_{\mathrm{B}}\sim 200\>\mathrm{mK}$ which only exceeds the lattice temperature during the QP burst by a factor of two. This is consistent with the observed decrease of the burst occurrence by about a factor of a few in the presence of gap difference.
@@
Conclusion.— In this work, we investigate the effect of superconducting gap engineering on QP bursts in transmon qubits. Our results confirm that the superconducting gap difference between the two sides of the junction exceeding the qubit energy, $\>\delta\Delta>hf_{q}$, can reduce both the burst detection rate, see Fig. 2(c), and the qubit relaxation rate during these events, see Fig. 2(e).
@@
However, we also uncover a critical limitation of this method: during bursts, the substrate and QPs retain high energy $\sim 90\>\mathrm{mK}$, only mildly lower than $\>\delta\Delta-hf_{q}$. We attribute this to the slow escape of hot phonons from our chip. Our findings suggest two potential strategies to further suppress the adverse effects of bursts. First, increasing the gap difference to $\>\delta\Delta/\>k_{\mathrm{B}}\gg 100\>\mathrm{mK}$ can prevent QPs from crossing the junction even when their energy is elevated. The increase of $\>\delta\Delta$ can be achieved with high-gap materials, such as granular aluminum [abeles_enhancement_1966, cohen_superconductivity_1968, deutscher_transition_1973, sun_measurements_2012]. Second, the effectiveness of gap engineering can be enhanced by providing an efficient escape mechanism for hot phonons. This involves improving the thermal link between the chip and the fridge’s base plate. Alternatively, the phonons can be absorbed by patches of normal metal placed directly on the chip [martinis_saving_2021, iaia_phonon_2022, larson_quasiparticle_2025]. Due to their high heat capacity, these patches can serve as efficient heat sinks. Combining these strategies should suppress the detrimental effect of QP bursts on superconducting quantum processors.

### v3

@@ Abstract
Using the common threshold for both devices, we see that the burst detection rate in the big-$\delta\Delta$ device ($0.57\>\mathrm{min}^{-1}$) is lower than that in the small-$\delta\Delta$ device ($2.76\>\mathrm{min}^{-1}$). Notably, the suppression of the burst detection rate by gap engineering is only about a factor of five. This observation is inconsistent with the QP energy thermalizing with the mixing chamber of the fridge immediately after the impact. If QPs were thermalized – as for resident QPs in Fig. 1 – $\delta\Delta$ would suppress the burst detection rate by about a factor of $\exp(\left[\>\delta\Delta-hf_{q}\right]/\>k_{\mathrm{B}}T)\sim 10^{4}$ in the big-$\delta\Delta$ device. Here, $T=25\>\mathrm{mK}$ is the base temperature of our fridge.
@@
Next, we consider how $T_{1}$ of the qubit returns to its steady state after the impact. To this end, we use continuous monitoring data from the measurement of Fig. 2(b-d). We identify the bursts using the threshold shown in Fig. 2(c). Averaging the measurement outcomes between many observed burst events allows us to determine the qubit relaxation rate $\Gamma_{10}$ as a function of time after the impact [noauthor_see_nodate]. The results of this analysis are shown in Fig. 2(e). For both devices, the return of the relaxation rate to the steady state is rougly exponential. The associated time constants $\sim 0.7\>\mathrm{ms}$ are insensitive to the gap engineering.
We note that different ionizing radiation sources, such as muons or gamma rays, may produce bursts of varying severity since vastly different amounts of energy can be absorbed by the chip [fowler_spectroscopic_2024]. Our measurements lack resolution to distinguish between different ionizing radiation sources. This is the reason why we average the response of all bursts together, regardless of their severity or origin. In the supplementary materials, we provide evidence that up to an overall rescaling, the dynamics of the qubit’s relaxation rate is independent of the burst severity [noauthor_see_nodate]. We leave a detailed investigation of different radiation sources to future work [vepsalainen_impact_2020, harrington_synchronous_2025, li_cosmic-ray-induced_2025, fowler_spectroscopic_2024].
[…]
@@
The qubit temperature $T_{q}$ surges to about 90 mK after the impact, significantly exceeding the fridge’s base temperature. The qubit temperature returns to its steady state exponentially, with a time constant of about $6\>\mathrm{ms}$. This time scale is almost 10 times longer than the $0.7\>\mathrm{ms}$ time constant for the decay of $x_{\mathrm{QP}}$. The temperature thus remains elevated even long after the generated by the impact non-equilibrium QPs are gone. This points to a dominant role of the substrate properties in determining the evolution of $T_{q}(t)$. Indeed, the observed increase in qubit temperature to $90\>\mathrm{mK}$ is consistent with the substrate phonon heat capacity for a typical radiation impact energy of $\sim 100\mathrm{keV}-1\>\mathrm{MeV}$ [fowler_spectroscopic_2024, noauthor_see_nodate]. The decay constant for the qubit temperature is determined by the rate of phonon escape from the system, likely through the chip clamps.
@@
The elevated substrate temperature after burst explains the weak sensitivity of burst occurrence to the gap difference. As previously described, the suppression of QP-induced qubit relaxation by $\delta\Delta$ is controlled by the activation exponent $\exp(-[\delta\Delta-hf_{q}]/\>k_{\mathrm{B}}T)$. For the gap difference to be effective, $\delta\Delta-hf_{q}$ must be much larger than $k_{\mathrm{B}}T$. For our big-$\delta\Delta$ device, $(\delta\Delta-hf_{q})/\>k_{\mathrm{B}}\sim 200\>\mathrm{mK}$ which only exceeds the lattice temperature during the QP burst by a factor of two. This is consistent with the observed five-fold decrease of the burst occurrence in the presence of gap difference.
At short times ($t\lesssim 100\>\mu\mathrm{s}$), $\Delta\Gamma_{10}(t)$ in the big-$\delta\Delta$ device deviates from exponential decrease, as shown in Fig. 2(e). We attribute this to the time dynamics of the QP distribution temperature. For $t\lesssim 100\>\mu\mathrm{s}$, the effective qubit temperature $T_{q}$ spikes well above 100 mK before rapidly decaying to $\sim 90$ mK for the remainder of the burst ($100\>\mu\mathrm{s}<t<5\>\mathrm{ms}$) [Fig. 2(e) inset and Fig. S14(f) in supplementary materials]. This temperature spike is likely related to the abundance of high-energy phonons that quickly get down-converted [kurilovich_correlated_2025]. Combined with the strong temperature dependence of $\gamma^{\mathrm{QP}}_{10}$ in big-$\delta\Delta$ device, the rapid spike and decay of the temperature is responsible for the non-exponential behavior of $\Delta\Gamma_{10}(t)$ during the initial part of the burst.
@@
Conclusion.— In this work, we investigate the effect of superconducting gap engineering on QP bursts in transmon qubits. Our results confirm that the superconducting gap difference between the two sides of the junction exceeding the qubit energy, $\delta\Delta>hf_{q}$, can reduce both the burst detection rate, see Fig. 2(c), and the qubit relaxation rate during these events, see Fig. 2(e).
@@
However, we also uncover a critical limitation of this method: during bursts, the substrate and QPs retain high energy $\sim 90\>\mathrm{mK}$, only mildly lower than $\delta\Delta-hf_{q}$. We attribute this to the slow escape of hot phonons from our chip. Our findings suggest two potential strategies to further suppress the adverse effects of bursts. First, increasing the gap difference to $\delta\Delta/\>k_{\mathrm{B}}\gg 100\>\mathrm{mK}$ can prevent QPs from crossing the junction even when their energy is elevated. The increase of $\delta\Delta$ can be achieved with high-gap materials, such as granular aluminum [abeles_enhancement_1966, cohen_superconductivity_1968, deutscher_transition_1973, sun_measurements_2012]. Second, the effectiveness of gap engineering can be enhanced by providing an efficient escape mechanism for hot phonons. This involves improving the thermal link between the chip and the fridge’s base plate. To achieve optimal thermalization, one should test various substrate materials and clamping schemes; this comparison is outside the scope of our work. Alternatively, the phonons can be absorbed by patches of normal metal placed directly on the chip [martinis_saving_2021, iaia_phonon_2022, larson_quasiparticle_2025]. Due to their high heat capacity, these patches can serve as efficient heat sinks. Combining these strategies should suppress the detrimental effect of QP bursts on superconducting quantum processors.